[figure: 9.7 cm]

IFLUENCE OF TEMPERATURE AND ANNEALING ON GMR

IN SPUTTERED PERMALLOY/Cu MULTILAYERS

M. Urbaniak^{a}, F. Stobiecki^{a}, T. Luciński^{a}, M. Kopcewicz^{b}, A. Grabias^{b} and

J. Aleksiejew^{a} aInstitute of Molecular Physics, Polish Academy of Sciences, Smoluchowskiego 17, PL 60-179 Poznań, Poland bInstitute of Electronics Materials Technology, Wólczyńska 133, 01-919 Warszawa, Poland

Influence of temperature and annealing on giant magnetoresistance (GMR) of Si(100)/ Cu(20nm)/Py(2nm)/(Cu(2nm)/Py(2nm))_{100} multilayer ( Py=Ni_{83}Fe_{17}) sputtered at room temperature in double face-to-face configuration is reported. It was found that GMR value, ΔR_{GMR}/R_{sat} (where R_{sat} is the resistance in saturation), monotonically decreases with increasing temperature (4.5% at 173K, about 1% at 373K). This results from the decrease of magnetic change of resistance, ΔR_{GMR}, and to the lesser extent from an increase of R_{sat}, though both of them are caused by the shortening of electrons mean free path. Observed almost linear decrease of GMR saturation field with increasing temperature is explained by temperature changes of magnetization profile. VSM measurements revealed that the increase in temperature results in pronounced decrease of remnant to saturation magnetisation ratio (M_{R}/M_{S}) suggesting that at low temperatures magnetic bridges between Py layers play an important role in magnetization process . It is shown that proper annealing, by an annihilation of bridges and/or lateral decoupling, leads to an increase of GMR ratio from 3.4% in as deposited state to 4.7%.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : M. Urbaniak

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences PL 60-179 Poznań Smoluchowskiego 17, Poland

Email address : urbaniak@ifmpan.poznan.pl